A dental student is conducting a study on the number of people who visit their dentist regularly. Of the 520 people surveyed, 312 indicated that they had visited their dentist within the past year.

Find the population proportion, as well as the mean and standard deviation of the sampling distribution for samples of size n=60.

Round all answers to 3 decimal places.

p=

Up=

Op=
Answer #1

population proportion p=312/520 =0.6000

for mean of  the sampling distribution phat=population proportion =0.60

std deviation of  the sampling distribution Se =(√(p*(1-p)/n))*(√((N-n)/(N-1))= 0.060